# Positive Semidefinite Matrix Supermartingales

### v4

@@ 3.2 Matrix Test Supermartingales
Scalar-valued nonnegative supermartingales are often constructed incrementally: if one can find an adapted nonnegative process $\{E_{n}\}$ that has the conditional expectation bound $\mathbb{E}(E_{n}|\mathcal{F}_{n})\leqslant 1$, then their product $Y_{n}=\prod_{i=1}^{n}E_{i}$ is a nonnegative supermartingale. In particular, the supermartingale satisfies $Y_{0}=1$ and thus satisfies (VI) with $\mathbb{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$. It is called a test supermartingale. We propose the following analogous way to construct $\mathcal{S}_{d}^{+}$-valued supermartingales that involves a bit more flexibility.
Lemma 3.5 (Construction of Matrix Test Supermartingales).
@@
with $\mathbf{Y}_{0}=\mathbf{I}$. It is called a matrix test supermartingale.
@@
With Lemma 3.5, we can recover matrix versions of several scalar test supermartingales in the recent sequential statistics literature. In the next subsection, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to two sequences of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 3.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (15). Combining with (MVI), we form the following generic scheme for sequentially testing a null hypothesis $\mathcal{H}_{0}$ using a PSD supermartingale outlined in Algorithm 1, where if $\mathcal{H}_{0}$ is true the procedure has at most $\alpha$ probability ever rejecting it.
@@ 3.3 Examples: Conditional Mean Testing Problems
We have numerous examples for Lemma 3.5 where one can conduct sequential hypothesis testing with Algorithm 1. All examples below, we note, involve the null hypothesis in the form of “the conditional mean of $\mathbf{X}_{n}$ equals $\mathbf{M}_{n}$” where $\{\mathbf{M}_{n}\}$ is a pre-specified predictable sequence. A common special case is when the matrices $\{\mathbf{X}_{n}\}$ are i.i.d. and one tests if their common mean equals a specific value.
@@
It is worth noting that in general it is hard to prove matrix MGF bounds and they are known only in very few cases. On the other hand, much more is known on bounding the matrix log-MGF $\log\left(\mathbf{G}_{n}(\cdot)\right)$ under various distribution assumptions, obtained via the Lieb-Tropp technique mentioned in Section 2.3.1. These log-MGF upper bounds only lead to scalar-valued supermartingales (9) instead of matrix-valued supermartingales. We review some existing results of matrix MGFs by Tropp [2012, 2015] in Section C.1.
@@
$-\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})}<\gamma_{n}<\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}$ (21)
@@
To see that this is a valid $(\mathbf{E}_{n},\mathbf{F}_{n})$ pair for Lemma 3.5, it is clear that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})=\mathbf{I}$ and (21) ensures $\mathbf{E}_{n}\succeq 0$.
@@
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$ under eight different conditions. These directly lead to $\mathcal{S}_{d}^{+}$-valued supermartingales via Lemma 3.5. However, Howard et al. [2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (15); they pursue scalar supermartingales instead, with methods we discuss later in Section 5.2. Of their eight inequalities, we will introduce two interesting cases now and (re-)prove them in Section B.2, omitting the other six due to their similar construction.
@@
Example 3.8 (Self-Normalized Matrix Supermartingale).
@@
Example 3.9.
@@
Example 3.10 (Empirical Bernstein-Type Matrix Supermartingale).
@@
An increasing function $\phi:\mathbb{R}\to\mathbb{R}$ is upper $p$-logarithmically contractive if $\phi(x)\leqslant\log(1+x+|x|^{p}/p)$; lower $p$-logarithmically contractive if $-\log(1-x+|x|^{p}/p)\leqslant\phi(x)$.
@@
Example 3.12 is proved in Section B.2. We remark that, in the scalar case, the two supermartingales constructed this way remain supermartingales under one-sided nulls, one for each side. This fails to hold with matrices (e.g., the null that $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{M}_{n}$) as there is no guarantee that $\exp(\pm\phi(\cdot))$ are operator monotone. Still, the two matrix test supermartingales play the role of detecting null-deviating behavior of opposite directions, and a powerful test for the point null $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ can be formed by a convex combination of the two.
@@ 3.4 Matrix e-Values and e-Processes
Throughout this subsection, we consider $X_{1},X_{2},\dots$ a sequence of random observations taking values in some space $\mathcal{X}$ and they generate the filtration $\{\mathcal{F}_{n}\}$. We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for $P$” (e.g. we may say an adapted stochastic process is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$, where $P$ is some distribution on $\mathcal{X}\times\mathcal{X}\times\dots$.
@@
and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}.$ (31)
@@
It can be seen from the matrix optional stopping theorem (Theorem 3.3) that if $\{\mathbf{Y}_{n}\}$ is a PSD supermartingale for $\mathcal{P}$ then it is also an e-process for $\mathcal{P}$. More generally, $\{\mathbf{Y}_{n}\}$ is a matrix e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{P}\}$ with $\mathbf{M}_{0}^{P}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{P}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $P$. Similar to the scalar case above, matrix e-processes satisfy (MVI), both time-uniform and at a stopping time:
@@
and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1}),$ (33)
@@ 4.1 Submartingales and Maximal Inequalities: From Forward to Backward
Substituting $M_{n}$ back to the inequality above, we have,
@@
The desired inequality is immediately proved by appealing to the fact that
$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.\qed$ (48)
@@ 4.2 Application: Concentration of Exchangeable Matrices
The scalar exchangeable Chebyshev inequality was recently proved by Ramdas and Manole [2024, Theorem 2]. Recall from Fact 1 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
@@
Let $p\in[1,2]$, and $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued, $L^{p}$ random matrices. Suppose $\mathbf{C}\in\mathcal{S}_{d}$ is deterministic, and either of the following holds:
@@
Recalling that the conditional expectation w.r.t. the exchangeable $\sigma$-algebra $\mathcal{E}_{n+1}$ equals the average over all $(n+1)!$ permutations of the indices $11,2,\dots,n+$ 1 (see e.g. Klenke [2013, Theorem 12.10]), we have
@@
(${{\succeq}}$ $\left(\frac{\sum_{\sigma\in\mathrm{S}_{n+1}}\left(\frac{\mathbf{X}_{\sigma(1)}+\dots+\mathbf{X}_{\sigma(n)}}{n}-\mathbf{C}\right)}{(n+1)!}\right)^{p}=\mathbf{R}_{n+1},$ (53)
@@
where the last step is due to Lemma B.2. ∎
@@
We remark that, from Theorem 4.5 to Theorem 4.6, we weaken the second moment to a $p$th moment assumption, but in an imperfect manner as we obtain a concentration not around the mean $\mathbf{M}$ but a Markov-type “upper tail bound” for $\mathcal{S}_{d}^{+}$-valued matrices instead. This is largely because the scalar map $x\mapsto|x|^{p}$ is convex on (,)($\infty$,$\infty$)( - ∞ , ∞ ) but not its matrix extension, and so the raw $p$th moment is used above. Also, Theorem 4.6 fails if $p>2$ because the scalar map $x\mapsto x^{p}$ is convex on $[0,\infty)$ but not its matrix extension. Nonetheless, thanks to Fact 3, we can resort to trace scalarization to obtain the following spectral and trace inequalities that hold for all $p\geqslant 1$.
@@
The inequality above, which we defer its straightforward proof to Section B.3, differs from Theorem 4.6 in that the concentration is now around the mean $\mathbf{M}$.

### v6

@@ 3.2 Matrix Test Supermartingales
Scalar-valued nonnegative supermartingales are often constructed incrementally: if one can find an adapted nonnegative process $\{E_{n}\}$ that has the conditional expectation bound $\mathbb{E}(E_{n}|\mathcal{F}_{n})\leqslant 1$, then their product $Y_{n}=\prod_{i=1}^{n}E_{i}$ is a nonnegative supermartingale which satisfies $Y_{0}=1$ and thus satisfies (VI) with $\mathbb{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$.
We propose the following analogous way to construct $\mathcal{S}_{d}^{+}$-valued supermartingales that involves a bit more flexibility.
Lemma 3.5 (Incremental Construction of Matrix Supermartingales).
@@
with $\mathbf{Y}_{0}=\mathbf{I}$.
@@
Recall from Section 2.2 the concept of testing a null hypothesis $\mathcal{H}_{0}$ with a scalar test supermartingale $\{Y_{n}\}$: a process with $Y_{0}=1$ that is a nonnegative supermartingale under $\mathcal{H}_{0}$. With Lemma 3.5, we can recover matrix versions of several scalar test supermartingales in the recent sequential statistics literature. In the next subsection, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to two sequences of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 3.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (15). Combining with (MVI), we form the following generic scheme for sequentially testing a null hypothesis $\mathcal{H}_{0}$ using a PSD supermartingale outlined in Algorithm 1, where if $\mathcal{H}_{0}$ is true the procedure has at most $\alpha$ probability ever rejecting it.
@@ 3.3 Examples: Conditional Mean Testing Problems
We have numerous examples for Lemma 3.5 where one can conduct sequential hypothesis testing with Algorithm 1. All examples below, we note, involve the null hypothesis in the form of “the conditional mean of $\mathbf{X}_{n}$ equals $\mathbf{M}_{n}$” where $\{\mathbf{M}_{n}\}$ is a pre-specified predictable sequence. A common special case is when the matrices $\{\mathbf{X}_{n}\}$ are i.i.d. and one tests if their common mean equals a specific value. It is not straightforward to see that some of these examples satisfy the assumptions of Lemma 3.5, so we prove them in Section 8.2.
@@
It is worth noting that in general it is hard to prove matrix MGF bounds and they are known only in very few cases. On the other hand, much more is known on bounding the matrix log-MGF $\log\left(\mathbf{G}_{n}(\cdot)\right)$ under various distribution assumptions, obtained via the Lieb-Tropp technique mentioned in Section 2.3.1. These log-MGF upper bounds only lead to scalar-valued supermartingales (9) instead of matrix-valued supermartingales. We review some existing results of matrix MGFs by Tropp [2012, 2015] in Section 9.1.
@@
$\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})}<\gamma_{n}<\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}$ (21)
@@
To see that this is a valid $(\mathbf{E}_{n},\mathbf{F}_{n})$ pair for Lemma 3.5, it is clear that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})=\mathbf{I}$ and (21) ensures $\mathbf{E}_{n}\succeq 0$. The resulting $\{\mathbf{Y}_{n}\}$ is therefore an $\mathcal{S}_{d}^{+}$-valued matrix martingale.
Remark 3.8.
Using Example 3.7 to test the null $\mathcal{H}_{0}:\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ with Algorithm 1, while any choice of predictable $\{\gamma_{n}\}$ that satisfies (21) makes sure the type-I error $\leqslant\alpha$ under $\mathcal{H}_{0}$, different choices of the sequence lead to difference in power when $\mathcal{H}_{0}$ does not hold. A simple example is stated below.
Suppose $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}+c\mathbf{I}$ for some unknown $c$ ($c=0$ under null $\mathcal{H}_{0}$). Since $\mathbb{E}[\mathbf{E}_{n}|\mathcal{F}_{n-1}]=(1+\gamma_{n}c)\mathbf{I}$, we can see from the proof of Lemma 3.5 that the “growth” event $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$ happens if and only if the predictable $\gamma_{n}$ and the ground-truth $c$ have matching signs: $\gamma_{n}c\geqslant 0$. If the sign of $c$ is not known beforehand, there are two ways to achieve this:
  • Since $\gamma_{n}$ can be picked predictably, one can estimate $c$ from $\mathbf{X}_{1},\dots,\mathbf{X}_{n-1}$ and match the sign of $\gamma_{n}$ with this estimator. For example, $\hat{c}_{n-1}=\mathrm{tr}((\mathbf{X}_{1}-\mathbf{M}_{1})+\cdots+(\mathbf{X}_{n-1}-\mathbf{M}_{n-1}))/(d(n-1))$, and $\gamma_{n}=\operatorname{sgn}(\hat{c}_{n-1})\gamma$ for some $0<\gamma<\min\left\{\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})},\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}\right\}$.
  • (Hedging). On the other hand, one may launch two test martingales, $\{\mathbf{Y}_{n}^{+}\}$ with $\gamma_{n}>0$, and $\{\mathbf{Y}_{n}^{-}\}$ with $\gamma_{n}<0$ for all $n$. Since one of them must match the correct sign of $c$ if $c\neq 0$, one can use $\{(\mathbf{Y}_{n}^{+}+\mathbf{Y}_{n}^{-})/2\}$ as the test process (which is still a test martingale for the null $\mathcal{H}_{0}:c=0$) to have power in detecting any $c\neq 0$.
The same logic applies to many of the subsequent examples.
@@
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$ under eight different conditions. These directly lead to $\mathcal{S}_{d}^{+}$-valued supermartingales via Lemma 3.5. However, Howard et al. [2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (15); they pursue scalar supermartingales instead, with methods we discuss later in Section 5.2. Of their eight inequalities, we will introduce two interesting cases now and (re-)prove them in Section 8.2, omitting the other six due to their similar construction.
@@
Example 3.9 (Finite-Variance Self-Normalized Matrix Supermartingale).
@@
As a bibliographical remark, we follow the usage of De la Pena et al. [2009], Bercu and Touati [2008, 2019], among others, in adopting the name “self-normalized”. These authors refer to a martingale $\{M_{n}\}$ on $\{\mathcal{F}_{n}\}$ as “self-normalized” if maximal inequalities of $\{Y_{n}\}$ are derived via the predictable quadratic variation $\sum\mathbb{E}((Y_{n}-Y_{n-1})^{2}|\mathcal{F}_{n-1})$ and the total quadratic variaiton $\sum(Y_{n}-Y_{n-1})^{2}$. The summands of these two quadratic variations correspond loosely to the $\mathbf{V}_{n}$ and $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}$ terms in (23). The next example further demonstrates this concept.
@@
Example 3.10 (Symmetric Self-Normalized Matrix Supermartingale).
@@
Example 3.11 (Empirical Bernstein Matrix Supermartingale).
@@
An increasing function $\phi:\mathbb{R}\to\mathbb{R}$ is upper $p$-logarithmically contractive if $\phi(x)\leqslant\log(1+x+|x|^{p}/p)$; lower $p$-logarithmically contractive if $\log(1-x+|x|^{p}/p)\leqslant\phi(x)$.
@@
Example 3.13 is proved in Section 8.2. We remark that, in the scalar case, the two supermartingales constructed this way remain supermartingales under one-sided nulls, one for each side. This fails to hold with matrices (e.g., the null that $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{M}_{n}$) as there is no guarantee that $\exp(\pm\phi(\cdot))$ are operator monotone. Still, the two matrix test supermartingales play the role of detecting null-deviating behavior of opposite directions, and a powerful test for the point null $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ can be formed by a convex combination of the two.
@@ 3.4 Matrix e-Values and e-Processes
Throughout this subsection, we consider $X_{1},X_{2},\dots$ a sequence of random observations taking values in some space $\mathcal{X}$ and they generate the filtration $\{\mathcal{F}_{n}\}$. We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for PP” (e.g. we may say an adapted stochastic process is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$, where $P$ is some distribution on $\mathcal{X}\times\mathcal{X}\times\dots$.
@@
$\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}.$ (31)
@@
It can be seen from the matrix optional stopping theorem (Theorem 3.3) that if $\{\mathbf{Y}_{n}\}$ is a PSD supermartingale for $\mathcal{P}$ then it is also an e-process for $\mathcal{P}$. More generally, the following is a sufficient condition for a matrix process to be an e-process:
Proposition 3.16.
Suppose for any $P\in\mathcal{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{P}\}$ for $P$ with $\mathbf{M}_{0}^{P}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{P}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $P$. Then, $\{\mathbf{Y}_{n}\}$ is a matrix e-process for $\mathcal{P}$.
This can be seen by simply noting that $\mathbb{E}_{P}\mathbf{Y}_{\tau}\preceq\mathbb{E}_{P}\mathbf{M}_{\tau}^{P}\preceq\mathbf{I}$ for any $P,\tau$. While the converse to Proposition 3.16 (that it is a necessary condition for e-processes as well) is true in the scalar case [Ramdas et al., 2020, Lemma 6], the argument fails to extend to matrices as the construction for scalars relies on the Snell envelope which we have yet to find a suitable generalization for matrices due to the non-existence of maximum and minimum under $\preceq$ within $\mathcal{S}_{d}$.
Similar to the scalar case above, matrix e-processes satisfy (MVI), both time-uniform and at a stopping time:
@@
$\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1}),$ (33)
@@ 4.1 Submartingales and Maximal Inequalities: From Forward to Backward
Substituting $\mathbf{Y}_{n}$ back to the inequality above, we have,
@@
The desired inequality follows from the monotone convergence theorem: the random variable $\mathbbmss{1}_{\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}}$ is the pointwise limit of the increasing sequence of random variables $\{\mathbbmss{1}_{\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}}\}_{N\geqslant 1}$. ∎
@@ 4.2 Application: Concentration of Exchangeable Matrices
The scalar exchangeable Chebyshev inequality was recently proved by Ramdas and Manole [2024, Theorem 2]. Recall from Fact 1 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $1\leqslant p<2$) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
@@
Let $p\in[1,2]$, and $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with $\mathbb{E}\|\mathbf{X}_{1}\|^{p}<\infty$. Suppose $\mathbf{C}\in\mathcal{S}_{d}$ is deterministic, and either of the following holds:
@@
Recalling that the conditional expectation w.r.t. the exchangeable $\sigma$-algebra $\mathcal{E}_{n+1}$ equals the average over all $(n+1)!$ permutations of the indices $1,2,\dots,n+1$ (see e.g. Klenke [2013, Theorem 12.10]), we have
@@
[…]
@@
where the last step is due to Lemma 8.2. ∎
@@
We remark that, from Theorem 4.5 to Theorem 4.6, we weaken the second moment to a $p$th moment assumption, but in an imperfect manner as we obtain a concentration not around the mean $\mathbf{M}$ but a Markov-type “upper tail bound” for $\mathcal{S}_{d}^{+}$-valued matrices instead. The raw $p$th moment is used above. Also, $p>2$ is not allowed. The reason is summarized in Table 1.
$p$ $x\mapsto x^{p}$, $x\geqslant 0$ $\mathbf{X}\mapsto\mathbf{X}^{p}$, $\mathbf{X}\succeq 0$ $\mapsto|x|^{p}, xx\in\mathbb{R}$ $\mathbf{X}\mapsto(\operatorname{abs}\mathbf{X})^{p}$, $\mathbf{X}\in\mathcal{S}_{d}$
$[1,2)$ convex convex convex non-convex
22 convex convex convex convex
$(2,\infty)$ convex non-convex convex non-convex
Table 1: Convexity of scalar-to-scalar vs. matrix-to-matrix power maps.
Nonetheless, thanks to Fact 3, we can resort to trace scalarization to obtain the following spectral and trace inequalities that hold for all $p\geqslant 1$.
@@
The inequality above, which we defer its straightforward proof to Section 8.3, differs from Theorem 4.6 in that the concentration is now around the mean $\mathbf{M}$.